# GSVA: Generalized Segmentation via Multimodal Large Language Models

### 15 Dec 2023 v1

@@ Abstract
Generalized Referring Expression Segmentation (GRES) extends the scope of classic RES to referring to multiple objects in one expression or identifying the empty targets absent in the image. GRES poses challenges in modeling the complex spatial relationships of the instances in the image and identifying non-existing referents. Recently, Multimodal Large Language Models (MLLMs) have shown tremendous progress in these complicated vision-language tasks. Connecting Large Language Models (LLMs) and vision models, MLLMs are proficient in understanding contexts with visual inputs. Among them, LISA, as a representative, adopts a special [SEG] token to prompt a segmentation mask decoder, e.g., SAM, to enable MLLMs in the RES task. However, existing solutions to of GRES remain unsatisfactory since current segmentation MLLMs cannot properly handle the cases where users might reference multiple subjects in a singular prompt or provide descriptions incongruent with any image target. In this paper, we propose Generalized Segmentation Vision Assistant (GSVA) to address this gap. Specifically, GSVA reuses the [SEG] token to prompt the segmentation model towards supporting multiple mask references simultaneously and innovatively learns to generate a [REJ] token to reject the null targets explicitly. Experiments validate GSVA’s efficacy in resolving the GRES issue, marking a notable enhancement and setting a new record on the GRES benchmark gRefCOCO dataset. GSVA also proves effective across various classic referring expression segmentation and comprehension tasks. Code will be available at https://github.com/LeapLabTHU/GSVA.
@@ 1 Introduction
Referring Expression Segmentation (RES) [7, 18] is an emerging vision-language (VL) task that aims to predict the masks of the corresponding objects referred to in the language expression. RES has great potential in many areas, especially embodied AI [57, 43, 16, 49], including VL navigation, human-robot interaction, etc. Nevertheless, the simplification in RES formulation that one referring expression must match an individual object in the image [22] has left a gap between current RES algorithms and real-world applications, neglecting multiple-target and empty-target cases.
@@
To bridge this gap, the Generalized Referring Expression Segmentation (GRES) [30] task has recently been proposed to support multiple-target and empty-target cases. In practical scenarios, users tend to refer to multiple subjects within a single instruction or provide descriptions that do not correspond to any targets in the image. As an example shown in Figure 1 (a), the left man and the second man from the right are targeted simultaneously, while a banana is referred to in a scene of apples in Figure 1 (b). RES takes no account of these cases, which are handled by GRES. In addition to the multimodal correspondences between images and text prompts in classic RES, GRES poses new challenges in handling more complicated multiple-target and empty-target cases. Therefore, the models are required to handle complex spatial relationships of the instances in the image [30] to segment the targets at various locations and reject the empty targets in the wrong places.
@@
The recent blooming Multimodal Large Language Models (MLLMs) [1, 68, 64, 31, 26] meet the requirements of GRES since they show excellency in complex reasoning [9] and instruction following [41] with visual inputs by aligning the LLMs [2, 40, 52, 53, 51, 8] and Visual Foundation Models (VFMs) [44, 54, 14, 60]. To support segmentation, LISA [24], a representative MLLM, incorporates LLaVA [31] and SAM [23] by a special [SEG] token as the prompt query, bridging the MLLM and the Segmentation Foundation Model (SFM) to generate masks of the referents in the user’s instructions. Successfully though LISA handles RES, GRES is still beyond its reach. As shown in Figure 1, LISA fails to work well in GRES where the multiple-target and empty-target challenges remain uncharted.
@@
To address the above challenges, we propose Generalized Segmentation Vision Assistant (GSVA). We attribute the vulnerability of LISA in GRES to the single permanent [SEG] token that restricts its flexibility. Therefore, we present two pivotal designs in GSVA: (1) learning to predict multiple [SEG] tokens to segment multiple targets; (2) rejecting empty targets in referring expressions by predicting [REJ] tokens. Specifically, when multiple targets are requested in the referring expression, we change the output sequence from a single [SEG] token to multiple [SEG] tokens corresponding to the entities in the expression, which encourages the MLLM to learn to output multiple [SEG] tokens. To distinguish each [SEG] token and avoid ambiguity, we add the expression of each entity in front of the corresponding [SEG] token, which can also be regarded as implicit In-Context Learning, hinting each [SEG] token to focus on the specific target. Meanwhile, if referents are absent in the image, the corresponding [SEG] tokens after the prompts are altered to [REJ] tokens to identify empty targets. The predicted [REJ] tokens are directly assigned with empty masks without being fed into the SFM, which liberates the segmentation model from seeking non-existing targets in the image. Benefiting from these novel designs, GSVA takes a big step forward in addressing GRES challenges, as shown in the second row of Figure 1.
In summary, our contributions are as follows: we propose two novel designs for segmentation MLLM to address the GRES challenges, including learning multiple [SEG] tokens along with the referring expressions and rejecting empty targets via [REJ] tokens. Extensive experiments demonstrate the effectiveness of the proposed GSVA. On the GRES benchmark gRefCOCO [30], GSVA outperforms LISA [24] by large margins, also surpassing previous SOTA methods, which sets new records with over 70% gIoU and cIoU. In addition, extensive experiments in RefCOCO, RefCOCO+ [22], and RefCOCOg [37] demonstrate that GSVA is also competitive on the classic RES and Referring Expression Comprehension (REC) tasks, achieving nearly 80% cIoU for RES and 90% Prec@0.5 for REC.
@@ 2 Related Works
RES and GRES. Referring Expression Segmentation (RES) [7, 18, 6, 37] aims to segment targets in the image based on language expressions. Under the assumption that one expression matches one existing target, many works explore fusing image and language [29, 27, 4, 15, 62]. Currently, most RES methods adopt the cross-attention module to bridge the modality gap [48, 63, 70, 59]. Another line of research enables the text prompts for segmentation model with a unified decoder architecture, including SAM [23], X-Decoder [75], SEEM [76], and Grounded-SAM which combines GroundingDINO [33] with SAM, etc. Although these works open a brilliant chapter for RES, the one-to-one restriction still holds. To break the jail for the arbitrary number of targets, DMMI [19] focuses on the one-to-many setting where text expression refers to varying numbers of targets. ReLA [30] proposes the Generalized Referring Expression Segmentation (GRES) task, supporting both the multi-target and empty-target scenarios, which is the main scope of our research of GSVA.
@@ 3.2 GRES: Task and Challenges
Differences from ReasonSeg. The Reasoning Segmentation (ReasonSeg) task is proposed by LISA [24], which emphasizes the complex text instructions in RES. In ReasonSeg, the instructions are more implicit and sophisticated than RES, including common sense reasoning and world knowledge. The logic chain is usually longer and more challenging in ReasonSeg, which requires the model to reason from the implicit query texts and deduce the target object referred to in the image. Nevertheless, GRES introduces complexity to RES in another dimension by involving complicated spatial relationships of objects. Hence, the model has to learn this spatial complexity and be capable of understanding the relationships between these instances. To meet these requirements, LISA tunes MLLMs with logically complex instructions paired with masks. In contrast, our proposed GSVA arouses the spatial modeling capabilities of MLLMs by learning multiple targets and rejecting empty targets. These designs are elaborated as follows.
@@ 3.3 Multiple [SEG] Tokens for Multiple Targets
User: What is {obj} in this image? Please output segmentation mask.
Assistant: Sure, it is [SEG].
@@ 3.4 Rejecting Empty Targets via [REJ] Tokens
User: What are {obj1} (does not exist), {obj2} in this image? Please output the segmentation masks. Assistant: Sure, {obj1}:[REJ], {obj2}:[SEG].
@@ 4.2 Referring Expression Segmentation
Settings. To validate the ablitiies to handle various tasks, we evaluate GSVA in the classic RES task. Following the common evaluation protocols, we test variants of GSVA and LISA equipped with different LLMs on RefCOCO, RefCOCO+ [22], and RefCOCOg [37]. We follow the UNC split to perform experiments on RefCOCO and RefCOCO+, and UMD split for RefCOCOg. The models are firstly pretrained as in GRES, and then finetuned for 10 epochs with a joint dataset of these three RES training set. The cIoU metric is adopted to measure the model performances.
@@ 4.3 Referring Expression Comprehension
Results. As shown in Table 3, we mainly compare our GSVA to LISA in the three variants. Without finetuning, GSVA-Vicuna-7B outperforms LISA with a large margin over 7% in almost all evaluation sets. The similar trends also hold for the Vicuna-13B and Llama2-13B variants. GSVA also benefits a lot from finetuning, e.g., the 7B variant achieves consistenly higher Prec@0.5 than uLLaVA-7B [61] with LoRA finetuning, which is another strong baseline that adopts the “mask2bbox” pipelines without direct bounding box supervision. The finetuned GSVA also shows competitive performances to the fully finetuned uLLaVA-7B, suggesting the strong potential of our method. With larger LLMs incorporated, the performances of GSVA continue with over 3% increments over all datasets.
@@ 4.4 Ablation Study
Answers without referring expression. To examine the efficacy of the addition of referring expression in front of [SEG] token, we remove all the referring expressions before the [SEG] tokens. Based on the removal of multiple [SEG]s and [REJ]s, erasing the added referring expression falls back to the original LISA model, as shown in ${}^{\text{th}}$ row of Table 4. To further study the of this addon, we choose only to remove it from GSVA model, keeping other configurations unchanged. Specifically, if there are two referents, the prompts in answer will turn to Assistant: Sure, [SEG], [SEG]. This modification leads to the last row in Table 4. The zero N-acc shows the model cannot learn to identify empty target without the help of the expressions, meanwhile the poor gIoU and cIoU indicates the segmentation ability is damaged since the ambiguity in the anonymous [SEG] tokens. This phenomenon also suggests that the added referring expression hint GSVA to associate each [SEG] token to its corresponding target, which is in coherence with our hypothesis.
@@ Acknowledgments
We thank Dr. Yuan Yao and Prof. Zhiyuan Liu for their insightful and valuable comments on this research project. We also appreciate their support of computational resources.
@@
Appendix A Implementation Details
@@ Appendix B GSVA on Ref-ZOM
Appendix C GSVA on Semantic Segmentation
@@
Appendix D More Ablation Study
@@
Weight sharing for different [SEG] tokens. Another design choice considering multiple [SEG] tokens is to use a set of individual [SEG] embeddings rather than sharing weights in GSVA. We adopt 8 tokens from [SEG000] to [SEG007] to study this design. In case of more than 8 targets, we dispatch the targets based on the remainder after divided by 8. However, this design poses convergence difficulties, as shown in Table 8. We attribute the intensive performance degradation to the slow convergence of the [SEG] token embedding, which is that too many [SEG] token embeddings could interfere with each other. Therefore it is reasonable to share weights between [SEG] tokens.
@@
Different SAM Backbones. We alter the default pretrained SAM-ViT-H [23] vision encoder to SAM-ViT-L and SAM-ViT-B, to verify the effectiveness of GSVA that is stable to the change of segmentation vision encoder $F_{\text{V2}}$. As shown in Table 9, when we substitute the ViT-H [13] backbones with SAM pretrained ViT-B/L, the segmentation performances decreases by 4$\sim$5while the accuracy of empty targets change mildly and even increases to 58.57% with ViT-B. The slight drops of gIoU and cIoU scores could be attributed to the usages of more miniature vision encoders, and the stable, even higher N-acc implies that generating [REJ] token by MLLM consistently works well with SFMs in different capacities. These results also demonstrate the effectiveness of the design that rejects the empty targets through a special token rather than producing an all-zero mask.
@@
Appendix E More Visualizations

### 21 Mar 2024 v3

@@ Abstract
Generalized Referring Expression Segmentation (GRES) extends the scope of classic RES to refer to multiple objects in one expression or identify the empty targets absent in the image. GRES poses challenges in modeling the complex spatial relationships of the instances in the image and identifying non-existing referents. Multimodal Large Language Models (MLLMs) have recently shown tremendous progress in these complicated vision-language tasks. Connecting Large Language Models (LLMs) and vision models, MLLMs are proficient in understanding contexts with visual inputs. Among them, LISA, as a representative, adopts a special [SEG] token to prompt a segmentation mask decoder, e.g., SAM, to enable MLLMs in the RES task. However, existing solutions to GRES remain unsatisfactory since current segmentation MLLMs cannot correctly handle the cases where users might reference multiple subjects in a singular prompt or provide descriptions incongruent with any image target. In this paper, we propose Generalized Segmentation Vision Assistant (GSVA) to address this gap. Specifically, GSVA reuses the [SEG] token to prompt the segmentation model towards supporting multiple mask references simultaneously and innovatively learns to generate a [REJ] token to reject the null targets explicitly. Experiments validate GSVA’s efficacy in resolving the GRES issue, marking a notable enhancement and setting a new record on the GRES benchmark gRefCOCO dataset. GSVA also proves effective across various classic referring segmentation and comprehension tasks. Code is available at https://github.com/LeapLabTHU/GSVA.
@@ 1 Introduction
Referring Expression Segmentation (RES) [7, 28] is an emerging vision-language (VL) task predicting the masks of the interested objects referred to in the language expression. RES has great potential in many areas, especially embodied AI [74, 60, 19, 66, 31], including VL navigation, human-robot interaction, etc. Nevertheless, the simplification in RES formulation that one referring expression must match an individual object in the image [34] has left a gap between current RES algorithms and real-world applications, neglecting multiple-target and empty-target cases.
@@
To bridge this gap, Generalized Referring Expression Segmentation (GRES) [42] has recently been proposed to support multiple-target and empty-target cases. Practically, users refer to multiple subjects within a single instruction or provide descriptions that do not correspond to any targets in the image. As an example shown in Figure 1 (a), the left man and the second man from the right are targeted simultaneously, while a banana is referred to in a scene of apples in Figure 1 (b). RES takes no account of these cases, which GRES handles. In addition to the multimodal correspondences between images and text prompts in classic RES, GRES poses new challenges in handling more complicated multiple-target and empty-target cases. Therefore, the models must handle complex spatial relationships of the instances in the image [42] to segment the targets at various locations and reject the empty targets in the wrong places.
@@
The recent blooming Multimodal Large Language Models (MLLMs) [1, 95, 91, 43, 38] meet the requirements of GRES since they show excellency in complex reasoning [10] and instruction following [55] with visual inputs by aligning the LLMs [2, 54, 69, 70, 68, 8] and Visual Foundation Models (VFMs) [61, 71, 17, 87] which are typically various Vision Transformers [16, 84, 86, 56, 57, 20, 25] to perceive image or video inputs. To support segmentation output, many works [94, 82, 83] link an MLLM (e.g., LLaVA [43]) and a segmentation model (e.g., SAM [35]) by prompting the decoder with special token embeddings (e.g., [SEG] in LISA [36]) to generate masks of the referents in the user’s instructions. Although these models manage to handle RES, GRES is still beyond their reach. As shown in Figure 1, LISA fails to work well in GRES where the multiple-target and empty-target challenges remain uncharted.
@@
To address the above challenges, we propose Generalized Segmentation Vision Assistant (GSVA). We attribute the vulnerability of other segmentation MLLMs in GRES to the single constant [SEG] token that restricts its flexibility. Therefore, we present two pivotal designs in GSVA: (1) learning to predict multiple [SEG] tokens to segment multiple targets; (2) rejecting empty targets in referring expressions by predicting [REJ] tokens. Specifically, when multiple targets are requested in the referring expression, we place multiple weight-sharing [SEG] tokens corresponding to the entities in the expression, encouraging the MLLM to learn to output multiple [SEG] tokens. To distinguish each [SEG] token and avoid ambiguity, we add the expression of each entity in front of the corresponding [SEG] token, hinting each [SEG] token to focus on the specific target, which can be regarded as implicit In-Context Learning, and dynamic neural network [21]. Meanwhile, if the referents are absent in the image, the corresponding [SEG] tokens after the prompts are altered to [REJ] tokens to identify empty targets. The predicted [REJ] tokens are directly assigned with empty masks without decoded, which liberates the segmentation model from seeking non-existing targets in the image. This Benefiting from these novel designs, GSVA takes a big step forward in addressing GRES challenges, as shown in the second row of Figure 1.
Our contributions are summarized as follows: (1) We propose GSVA to solve the GRES problem with MLLM by handling the spatial relationships among targets, and study the GRES problem systematically in the context of LLM for the first time. (2) We propose the non-trivial shared-weight multiple [SEG] tokens guided by each referent prompt to address the multiple-target problem. (3) We firstly propose a clean solution, the [REJ] token, to reject the empty targets, which can be seamlessly applied to various models. (4) GSVA is intuitive and effective, achieving state-of-the-art performance on the GRES benchmarks.
@@ 2 Related Works
RES and GRES. Referring Expression Segmentation (RES) [7, 28, 6, 49] assumes that one expression matches one existing target, and many works explore fusing image and language [41, 39, 4, 18, 89] to segment objects under instructions. Currently, most RES methods adopt the cross-attention module or cross-modal alignment to bridge the modality gap [65, 90, 97, 81, 9]. Another line of research enables the text prompts for segmentation model with a unified decoder [35, 102, 103, 45], offering more flexible outputs. To break the jail for the arbitrary number of targets, DMMI [29] focuses on the one-to-many setting where text expression refers to varying numbers of targets.ReLA [42] proposes the Generalized Referring Expression Segmentation (GRES) task, supporting both the multi-target and empty-target scenarios, which is our main research scope.
@@
Dynamic Networks. Dynamic neural networks [21] can adapt their architectures [30, 76, 75, 77, 78, 23, 26, 22, 24] or parameters [59, 13, 14] to different inputs or switch the computation architecture in adjustment to different different time steps [85, 79, 52, 53], in order to achieve better accuracy and efficiency. In GSVA, the weight-sharing [SEG] tokens adapt to multiple targets under the hint of the prepended target prompts and dynamically reject empty targets with an individual prediction of rejection tokens.
@@ 3.2 GRES: Task and Challenges
Differences from ReasonSeg. Reasoning Segmentation (ReasonSeg) proposed by LISA [36] emphasizes the complex text instructions in RES. In ReasonSeg, the instructions are more implicit and sophisticated, forcing models to reason using world knowledge. Besides, the logic chain is usually longer and more challenging in ReasonSeg, which requires the model to deduce the final target object referred to in the image. In contrast, GRES increases complexity in another dimension by involving complicated spatial relationships. Hence, the model has to learn to handle this spatial information and understand the relationships between the instances. To meet these requirements, LISA tunes MLLMs with complex instructions paired with masks, while GSVA arouses the spatial modeling capabilities of MLLMs by learning multiple targets and rejecting empty targets.
@@ 3.3 Multiple [SEG] Tokens for Multiple Targets
User: What is {obj} in this image? Please output segmentation mask. Assistant: Sure, it is [SEG].
@@ 3.4 Rejecting Empty Targets via [REJ] Tokens
User: What are {obj1} (absent), {obj2} (absent), {obj3} in this image? Please output the segmentation masks. Assistant: Sure, {obj1}:[REJ], {obj2}:[REJ], {obj3}:[SEG].
@@ 4.2 Referring Expression Segmentation
Settings. To validate the abilities to handle various tasks, we evaluate GSVA in the classic RES task. Following the common evaluation protocols, we test variants of GSVA and LISA equipped with different LLMs on RefCOCO, RefCOCO+ [34], and RefCOCOg [49]. We follow the UNC split to perform experiments on RefCOCO and RefCOCO+, and UMD split for RefCOCOg. The models are firstly pretrained as in GRES, and then finetuned for 10 epochs with a joint dataset of these three RES training set. The cIoU metric is adopted to measure the model performances.
@@ 4.3 Referring Expression Comprehension
Results. As shown in Table 3, we mainly compare our GSVA to LISA in the three variants. Without finetuning, GSVA-Vicuna-7B outperforms LISA with a large margin over 7% in almost all evaluation sets. The similar trends also hold for the Vicuna-13B and Llama2-13B variants. GSVA also benefits a lot from finetuning, e.g., the 7B variant achieves consistently higher Prec@0.5 than uLLaVA-7B [88] with LoRA finetuning, which is another strong baseline that adopts the “mask2bbox” pipelines without direct bounding box supervision. The finetuned GSVA also shows competitive performances to the fully finetuned uLLaVA-7B, suggesting the strong potential of our method. With larger LLMs incorporated, the performances of GSVA continue with over 3% increments over all datasets.
@@ 4.4 Ablation Study
Answers without referring expression. To examine the efficacy of the hinting prompts, we remove all the referring expressions before the [SEG] tokens. Based on the removal of multiple [SEG]s and [REJ]s, erasing the added referring expression falls back to the original LISA model, as shown in ${}^{\text{th}}$ row. We further choose only to remove it from GSVA model, keeping other configurations unchanged. Specifically, if there are two referents, the prompts in the answer will turn to Assistant: Sure, [SEG], [SEG]., whose results are in the last row. The zero N-acc shows the model fails to identify any empty target without the help of the expressions, meanwhile the poor gIoU and cIoU indicates the segmentation ability is damaged. This phenomenon also suggests that the added referring expression hint GSVA to associate each [SEG] token to its corresponding target, which is in coherence with our hypothesis.
@@ Acknowledgments
This work is supported in part by the National Key R&D Program of China under Grant 2021ZD0140407, the National Natural Science Foundation of China under Grants 62321005 and 62276150. We also thank Dr. Yuan Yao and Prof. Zhiyuan Liu for their insightful and valuable comments on this research project. We also appreciate their generous support of computational resources.
@@
Appendix A Discussions
A.1 Comparison to Concurrent Works
[…]
A.2 GSVA and Reasoning Segmentation
Method gIoU cIoU
LISA-7B (ft) 50.5 53.2
GSVA-7B (ft) 50.5 56.4
Table 5: Results of LISA and GSVA on ReasonSeg dataset.
We evaluate GSVA on ReasonSeg [36] with a generalized configuration. Table 5 demonstrates the capability of GSVA to handle instructions with complex logic, showing competitive results to LISA [^1] Results reproduced by the open-source code of LISA.. Figure 7 exemplifies that GSVA can segment the ReasonSeg referent while rejecting the additional empty target, while LISA fails to make the rejection, which further verifies the efficacy of GSVA.
A.3 Support of Various Question Types
Question gIoU cIoU N-acc.
“What” 63.32 61.70 56.44
“Where” 63.43 61.57 56.98
“Show” 63.05 61.35 56.48
“Outline” 63.16 61.51 56.27
Table 6: GRES results with different types of questions.
For simplicity, we show one type of question in the example in the paper. In contrast, the prompt types are diversified during training, including “Please segment {objs} in this image”, ”Can you segment {objs} in the image”, etc. Besides, the training data also contains VQA and Reasoning Segmentation, which include various questions and answers. As shown in Table 6, apart from the substituted “Where” question of “What,” we also examine the robustness of GSVA by testing the pretrained model with the unseen “Show” and “Outline” questions, short for “Show me {objs} in the image with segmentation masks” and “Outline {objs} in this image with segmentation masks,” respectively. As demonstrated in Figure 7, different question forms do not impact much performance deviation as long as they are reasonable.
Figure 5: Generalized Reasoning Segmentation Example.
Figure 6: Example of different types of questions.
Figure 5: Generalized Reasoning Segmentation Example.
Figure 6: Example of different types of questions.
Figure 7: Failure case example of GSVA in terms of N-acc., where empty targets are incorrectly predicted with [SEG] tokens.
A.4 Multiple Objects in One Expression
In GSVA, each expression is separated with a comma in the prompt, corresponding to a segmentation map or rejection token. If more than one object is stated in the expression, a single [SEG] token will guide a segmentation mask to cover all objects. The common practice of GRES is merging all masks of the objects referred to into one as the ground truth for evaluation, which we follow for fair comparisons. Since there exist numerous expressions containing both present and absent objects, GSVA learns to predict the mask of the union of the referents so the absent objects will not occur.
A.5 Failure Cases on N-acc.
Although GSVA achieves a state-of-the-art level of N-acc on the GRES task and outperforms both Non-LLM and LLM models, we still observe that several failure cases are relatively small and vague fragments in the images, easily leading to misperception. As shown in Figure 7, there is no fridge obviously at the top left, but the model proposes a nearer fridge instead. It is hard for the model to tell whether in the corner stands a fridge, especially at low input resolution. This suggests using higher-resolution vision encoders, which will be an interesting future direction.
Appendix B Implementation Details
@@ Appendix C GSVA on Ref-ZOM
Appendix D GSVA on Semantic Segmentation
@@
Appendix E More Ablation Study
@@
Weight sharing for different [SEG] tokens. Another design choice considering multiple [SEG] tokens is to use a set of individual [SEG] embeddings rather than sharing weights in GSVA. We adopt 8 tokens from [SEG000] to [SEG007] to study this design. In case of more than 8 targets, we dispatch the targets based on the remainder after divided by 8. However, this design poses convergence difficulties, as shown in Table 10. We attribute the intensive performance degradation to the slow convergence of the [SEG] token embedding, which is that too many [SEG] token embeddings could interfere with each other. Therefore, it is reasonable to share weights between [SEG] tokens.
@@
Different SAM Backbones. We alter the default pretrained SAM-ViT-H [35] vision encoder to SAM-ViT-L and SAM-ViT-B, to verify the effectiveness of GSVA that is stable to the change of segmentation vision encoder $F_{\text{V2}}$. As shown in Table 11, when we substitute the ViT-H [16] backbones with SAM pretrained ViT-B/L, the segmentation performances decreases by 4$\sim$5while the accuracy of empty targets change mildly and even increases to 58.57% with ViT-B. The slight drops in gIoU and cIoU scores could be attributed to the usages of more miniature vision encoders, and the stable, even higher N-acc implies that generating [REJ] token by MLLM consistently works well with SFMs in different capacities. These results also demonstrate the effectiveness of the design that rejects the empty targets by a special token rather than predicting an all-zero mask.
@@
Appendix F More Visualizations